OpenStudy (anonymous):

plz help T~T how do I factor x^4-1

OpenStudy (solomonzelman):

well, at first use the difference of squares. \(\large\color{black}{ \displaystyle \left(x^2\right)^2-1^2}\)

OpenStudy (solomonzelman):

The rule (difference of squares) is: \(\large\color{black}{ \displaystyle a^2-b^2=(a-b)(a+b) }\)

OpenStudy (jdoe0001):

\(\bf \textit{differences of squares} \\ \quad \\ (a-b)(a+b) = a^2-b^2\qquad \qquad a^2-b^2 = (a-b)(a+b)\)

OpenStudy (anonymous):

oooohhhhhh thanks i know how to do the rest CX thank you

OpenStudy (solomonzelman):

ok, when you are done, blurt out the answer (jk) ... (just for verification).